OpenStudy (anonymous):

Help me with average rate of change calculations it's f(x+h)-f(x)/h and f(x)=x^4 with point (-1,1). How does it equal -4? I can't get it, I only get 4

OpenStudy (campbell_st):

ok... well this is differentiation b 1st principles... which will be long and tedious \[f(x + h) = (x + h)^4 = x^4 + 4x^3h + 6x^2h^2 + 4xh^3 + h^4 \] and \[f(x) = x^4\] substitute them and you should also have a limit \[\lim_{h \rightarrow 0}\frac{x^4 + 4x^3h + 6x^2h^2 + 4xh^3 + h^4 - x^4}{h}\] collect like terms in the numerator then factor the numerator look for common factors between the numerator and denominator substitute the limit and you will have f'(x) then find f'(-1), that is substitute x = -1 from the point and evaluate

OpenStudy (anonymous):

yeah I was dumb I didn't foil properly thanks lol